which conic section is 3(x+5)^2+3y-15=0
Answer by lwsshak3(11628)

which conic section is
3(x+5)^2+3y-15=0
3(x^2+10x+25)+3y-15=0
3x^2+30x+75+3y-15=0
3x^2+30x+3y=-60
x^2+10x+y=-20
complete the square:
(x^2+10x+25)+y=-20+25
(x+5)^2+y=5
y=-(x+5)^2+5
This is a parabola that opens down with vertex at (-5,5)